# FSR-VLN: Fast and Slow Reasoning for Vision-Language Navigation with Hierarchical Multi-modal Scene Graph

### v1

@@ I INTRODUCTION
    Inspired by the dual-process theory of human cognition, we introduce a novel Fast-to-Slow Navigation Reasoning system. The system operates in two distinct stages: a fast, intuitive matching step retrieves candidate views and objects, followed by a slow, deliberate reasoning stage where an LLM/VLM verifies and refines the final goal. This multi-stage architecture seamlessly integrates efficient feature-space matching with robust LLM/VLM-driven visual verification, leading to a significant improvement in success rates for real-world, long-range navigation.
@@ IV-A Experiment Setup
Dataset. To validate our system in the real world, we use a Unitree-G1 humanoid robot equipped with a calibrated Intel RealSense D455 RGBD camera and a Mid360 LiDAR (Figure 1) to collect LiDAR–camera data across long-range office environments with long corridors and multiple rooms, denoted as Room1, Room2, Room3, and Room4.
@@ IV-B Benchmark Results
TABLE I: Comparison of Methods with SR, RSR@Top1, and RSR@Top5 metrics. The metrics are visualized green (best) to red (worst).
Method Map Representation Time SR RSR@Top1 RSR@Top5
@@
TABLE II: Comparison of RSR@Top1 Performance Across Different FSR-VLN Settings. The highest value is highlighted in bold.
@@ IV-C Ablations Analysis
To further validate the effectiveness of HMSG representation and FSR, we examine the impact of spatial instructions and the FSR process on FSR-VLN performance, as summarized in Table II. Without Navigation Reasoning (NR) and Spatial Target (ST) instructions guide long-range navigation is guided by restricting object search to the target room, reducing global matching errors, and improving RSR. In long-range environments, this room-level guidance is particularly critical; when combined with the hierarchical scene graph, the search is further confined to the designated room, enhancing navigation success. With NR, extended VLM reasoning allows FSR to verify the correctness of fast matching, and VLM-based selection refinement further increases RSR, demonstrating the effectiveness of the approach.

### v2

@@ I INTRODUCTION
    Inspired by the dual-process theory of human cognition, we introduce a novel Fast-to-Slow Navigation Reasoning system. The system operates in two distinct stages: a fast, intuitive matching step retrieves candidate views and objects, followed by a slow, deliberate reasoning stage where an VLM verifies and refines the final goal. This multi-stage architecture seamlessly integrates efficient feature-space matching with robust VLM-driven visual verification, leading to a significant improvement in success rates for real-world, long-range navigation.
@@ IV-A Experiment Setup
Dataset. To validate our system in the real world, we use a Unitree-G1 humanoid robot equipped with a calibrated Intel RealSense D455 RGBD camera and a Mid360 LiDAR (Figure 1) to collect LiDAR–camera data across long-range office environments with long corridors and multiple rooms, denoted as Room1, Room2, Room3, and Room4. We also evaluated our approach on eight scenes from the HM3D-SEM dataset [49].
@@
In the HM3D-SEM dataset, as osmAG-LLM, we select HOV-SG and osmAG-LLM as our baseline because they share fundamental design principles with our approach[50].
@@ IV-B Benchmark Results
TABLE I: Object Retrieval Comparison with SR, RSR@Top1, and RSR@Top5 metrics. The metrics are visualized green (best) to red (worst).
Method Map Representation&Navigation Time SR RSR@Top1 RSR@Top5
@@
TABLE II: Object Retrieval Comparison with RSR@Top1 and RSR@Top5 on HM3DSEM. The metrics are visualized green (best) to red (worst).
Method Time RSR@Top1 RSR@Top5
1m 2m 3m 1m 2m 3m
HOVSG[20] 0.2s 0.52 0.64 0.70 0.76 0.82 0.88
osmAG-LLM[49] 3.0s 0.28 0.50 0.69 0.47 0.83 0.90
FSR-VLN (Ours) 5.5s 0.87 0.88 0.88 0.92 0.92 0.92
As illustrated in Table II, the Top-1 retrieval success rate of osmAG-LLM is significantly lower than that of HOVSG. The main reason is that osmAG-LLM loses the hierarchically rich visual open-vocabulary feature representation present in HOVSG and instead relies solely on textual recognition results. Although the osmAG-LLM semantic map is also hierarchical, it only preserves XML-level textual information without retaining visual CLIP features. In contrast, our method not only preserves CLIP-based visual embeddings but also further interacts with the original image information through VLM, leading to a notable improvement in retrieval success.
@@
TABLE III: Comparison of RSR@Top1 Performance Across Different FSR-VLN Settings. The highest value is highlighted in bold.
@@ IV-C Ablations Analysis
To further validate the effectiveness of HMSG representation and FSR, we examine the impact of spatial instructions and the FSR process on FSR-VLN performance, as summarized in Table III. Without Navigation Reasoning (NR) and Spatial Target (ST) instructions guide long-range navigation is guided by restricting object search to the target room, reducing global matching errors, and improving RSR. In long-range environments, this room-level guidance is particularly critical; when combined with the hierarchical scene graph, the search is further confined to the designated room, enhancing navigation success. With NR, extended VLM reasoning allows FSR to verify the correctness of fast matching, and VLM-based selection refinement further increases RSR, demonstrating the effectiveness of the approach.